Consider a firm with production function Q(L,K) = min( 6L , 3K ).
The wage rate is w = 6, and the rental rate of capital is r = 3.
(a) Calculate the optimal capital to labour ratio
(b) Derive the long-run cost function of the firm.

need help in all! Thank you!
1
Expert's answer
2020-12-08T07:48:32-0500

 Q(L,K) = min( 6L , 3K ).

Wage rate w = 6 rental rate r=3

a) optimal combination of labor and capital is given by 6L = 3K

K= 2L

"\frac{K}{L} = 2"

Optimal capital labor ratio = "(\frac{K}{L} = 2)"

b)Substitute the ratio K= 2L in Q

Q = min{6L,3(2L)} = min{6L,6L}

Q= 6L

"L= \frac{Q}{6}"

"K= \frac{2Q}{6}" = "\frac{Q}{3}"

Cost(C) = WL + rK

C= "\frac{Q}{6}(W)" + "\frac{Q}{3}(r)"

C="\frac{Q}{6}(6)" + "\frac{Q}{3}(3)"

C= 2Q

therefore long run cost function C(Q)

C(Q) = 2Q